# latexOverwriting file

### v1

@@ Abstract
For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level. In order to transpose convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics. In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling. Our approach extends Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems.
@@ 1 Introduction
This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants. The former, initially proposed in [46], is an algorithm for global optimization tasks, while the latter, proposed in [11], can be employed for both sampling and optimization purposes. Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels. This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
$\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in,$ (1.1)
where $J>1$ is a natural number, :={$:=\{1,\dots,J\}$ and $\mu^{J}_{t}:=\frac{1}{J}\sum_{i=1}^{J}\delta_{X^{(i)}_{t}}$ denotes the empirical measure associated with the particle system. For a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$, the notation $\mathcal{M}_{\beta}(\mu)$ denotes the mean of $\mu$ after reweighting by $\operatorname{e}^{-\beta f}$. Its precise definition, together with its expression when $\mu=\mu^{J}_{t}$, is given by
@@
The weighted mean $\mathcal{M}_{\beta}(\mu^{J}_{t})$ may be viewed as an approximation of the particle position $X^{(j)}_{t}$ such that $f(X^{(j)}_{t})$ is minimal. In (1.1), the processes $\bigl{(}W^{(j)}_{t}\bigr{)}_{j\in}$ are independent $\mathbf{R}^{d}$-valued Brownian motions and $S\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ is a globally Lipschitz continuous function. The two diffusion types commonly studied in the literature are isotropic diffusion [46, 10, 23], with $S(u)=\sqrt{2\theta}\lvert u\rvert\operatorname{I}_{d}$, and anisotropic or component-wise [13, 24] diffusion with $S(u)=\sqrt{2\theta}\operatorname{diag}(u)$. Here $\operatorname{I}_{d}\in\mathbf{R}^{d\times d}$ is the identity matrix and $\operatorname{diag}\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ the operator mapping a vector onto a diagonal matrix with the vector as its diagonal. It was shown in [13] that CBO with anisotropic diffusion is well-suited for high-dimensional global optimization problems, for example those arising in machine learning. Commonly a constant coefficient is included in front of the drift term in (1.1), but since this coefficient can be removed by rescaling time upon redefining the diffusion parameter $\theta$, we choose to omit it.
By design, CBO allows passage to the corresponding mean-field description. Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [37], it is possible to formally write down the corresponding mean-field equation:
@@
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [29]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [34, 10, 23]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [10]. The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$. The proof in [10] covers also non-convex cost functions $f$.
@@
The convergence, in an appropriate sense, of the empirical measure associated with the CBO finite-size particle system (1.1) to the solution of the mean-field equation (1.3) in the large particle limit $J\to\infty$ was first proved rigorously in a recent paper [34] using a compactness argument, as we discuss further in Subsection 1.3. However, the method of proof employed in [34] does not provide an explicit rate of convergence with respect to the number of particles. Obtaining such an explicit rate of convergence is highly desirable, as it would enable to transfer convergence results concerning the mean-field equation (1.3) to the finite ensemble size setting (1.1) by means of a triangle inequality argument, modulo appropriate interfacing of the metrics. This is the approach proposed in [23, Theorem 13], where the authors obtain a partial mean-field result in a set of high probability, enabling them to prove convergence with high probability to the global minimizer in the finite ensemble size setting. Here, we provide a stronger version of the mean-field limit result given in [23, Proposition 15], extending to more general diffusion terms and providing scope for global convergence results in a stronger metric than convergence in probability.
@@
A number of extensions of CBO have been proposed in the literature. As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [13]. Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [30, 31, 41]. (On this subject, we point out that the results in [31] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [35], memory effects [49], and Lévy noise [36], as well as extensions to multi-objective optimization [40, 7] and global optimization with constraints [21, 22, 12, 8, 32, 4]. Moreover, a variant tailored for objective functions with several minima was proposed in [9], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [29, 17]. Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [48].
@@ 1.2 Consensus-Based Sampling (CBS)
Consensus-Based Sampling is a sampling method inspired by the Ensemble Kalman Sampler (EKS) [26] and by consensus-based approaches such as CBO. It is based on an interacting particle system with a noise term preconditioned by the square root of the empirical covariance. The method in the finite ensemble size and continuous time setting reads
$\mathrm{d}X^{(j)}_{t}=-\Bigl{(}X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\,\mathrm{d}W^{(j)}_{t},\qquad j\in,$ (1.4)
@@
More precisely, the authors of [11] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$. In addition, it was noticed in [11] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [46]. Since the convergence analysis in [11] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$. This has so far been left as an open problem. Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [9], and for rare event estimation in [2].
@@ 1.3 Propagation of chaos
Although the consensus-based dynamics can be written in the form (1.6), it is not possible to do so in such a way that $\widetilde{b},\widetilde{\sigma},K_{b}$ and $K_{\sigma}$ are all globally Lipschitz continuous and at the same time $K_{b}$ and $K_{\sigma}$ are globally bounded. Therefore, the classical McKean’s theorem [15, Theorem 3.1] is not directly applicable, and a different approach is needed. We briefly discuss hereafter the existing mean-field results for CBO, noting that no results have been obtained so far for CBS.
  • In [34], propagation of chaos for CBO is proved through a compactness argument based on Prokhorov’s theorem. However, this method comes without convergence rates in terms of the number of particles $J$.
  • Later, propagation of chaos by means of synchronous coupling was shown with optimal rates in [21] for a variant of CBO constrained to hypersurfaces.
  • Recent results in [23] also show a probabilistic mean-field approximation of the form
@@
    It is additionally proved in [23], by application of Markov’s inequality, that the set of realizations $\Omega_{M}$ has probability close to 0 for large $M$. Specifically, it holds that $\mathbf{P}(\Omega_{M})=\mathcal{O}(M^{-1})$.
  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [36]. In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1. There are, however, several differences between our results and those presented in [36]. Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [1, 18] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3). Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [36].
@@ 1.4 Our contributions
  • We present two approaches for dealing with the weighted mean (in the case of CBO) and the weighted covariance (in the case of CBS). One of these approaches is based on the careful use of indicator sets, whereas the other is based on stopping times. The latter approach is inspired by a method initially proposed in [33] in order to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
@@ 2.1 Main assumptions
    for some constants $L_{f}>0$ and $s\geqslant-1$.
@@
    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
    $\forall x\in\mathbf{R}^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
@@
  • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [10]. The well-posedness and mean-field results proved in this paper generalize beyond this setting, as explained below.

### v2

# Mean-field limits for Consensus-Based Optimization and Sampling

@@ Abstract
For algorithms based on interacting particle systems that admit a mean-field description, convergence analysis is often more accessible at the mean-field level. In order to transpose convergence results obtained at the mean-field level to the finite ensemble size setting, it is desirable to show that the particle dynamics converge in an appropriate sense to the corresponding mean-field dynamics. In this paper, we prove quantitative mean-field limit results for two related interacting particle systems: Consensus-Based Optimization and Consensus-Based Sampling. Our approach requires a generalization of Sznitman’s classical argument: in order to circumvent issues related to the lack of global Lipschitz continuity of the coefficients, we discard an event of small probability, the contribution of which is controlled using moment estimates for the particle systems. In addition, we present new results on the well-posedness of the particle systems and their mean-field limit, and provide novel stability estimates for the weighted mean and the weighted covariance.
@@ 1 Introduction
This work focuses on the rigorous derivation of quantitative mean-field limits for recently proposed algorithms known as Consensus-Based Optimization (CBO) and Consensus-Based Sampling (CBS), as well as their variants. The former, initially proposed in [CBO], is an algorithm for global optimization tasks, while the latter, proposed in [CBS-Carrillo2021], can be employed for both sampling and optimization purposes. Both methods are based on an interacting particle system with only locally, not globally, Lipschitz coefficients and interaction kernels. This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
$\mathrm{d}X^{(j)}_{t}$ $=-\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}t+S\left(X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\right)\,\mathrm{d}W^{(j)}_{t},\qquad t\geqslant 0,\qquad j\in\llbracket 1,J\rrbracket,$ (1.1)
where $J>1$ is a natural number, $\llbracket 1,J\rrbracket:=\{1,\dots,J\}$ and $\mu^{J}_{t}:=\frac{1}{J}\sum_{i=1}^{J}\delta_{X^{(i)}_{t}}$ denotes the empirical measure associated with the particle system. For a probability measure $\mu\in\mathcal{P}(\mathbf{R}^{d})$, the notation $\mathcal{M}_{\beta}(\mu)$ denotes the mean of $\mu$ after reweighting by $\operatorname{e}^{-\beta f}$. Its precise definition, together with its expression when $\mu=\mu^{J}_{t}$, is given by
@@
The weighted mean $\mathcal{M}_{\beta}(\mu^{J}_{t})$ may be viewed as an approximation of the particle position $X^{(j)}_{t}$ such that $f(X^{(j)}_{t})$ is minimal. In (1.1), the processes $\bigl{(}W^{(j)}_{t}\bigr{)}_{j\in\llbracket 1,J\rrbracket}$ are independent $\mathbf{R}^{d}$-valued Brownian motions and $S\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ is a globally Lipschitz continuous function. The two diffusion types commonly studied in the literature are isotropic diffusion [CBO, carrillo2018analytical, fornasier2021consensusbased], with $S(u)=\sqrt{2\theta}\lvert u\rvert\operatorname{I}_{d}$, and anisotropic or component-wise [CBO-high-dim, FornasierKlockRiedl2022-anisotropic] diffusion with $S(u)=\sqrt{2\theta}\operatorname{diag}(u)$. Here $\operatorname{I}_{d}\in\mathbf{R}^{d\times d}$ is the identity matrix and $\operatorname{diag}\colon\mathbf{R}^{d}\to\mathbf{R}^{d\times d}$ the operator mapping a vector onto a diagonal matrix with the vector as its diagonal. It was shown in [CBO-high-dim] that CBO with anisotropic diffusion is well-suited for high-dimensional global optimization problems, for example those arising in machine learning. Commonly a constant coefficient is included in front of the drift term in (1.1), but since this coefficient can be removed by rescaling time upon redefining the diffusion parameter $\theta$, we choose to omit it.
By design, CBO allows passage to the corresponding mean-field description. Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [PSO], it is possible to formally write down the corresponding mean-field equation:
@@
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical]. The distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large exponential parameter $\beta$. The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@
The convergence, in an appropriate sense, of the empirical measure associated with the CBO finite-size particle system (1.1) to the solution of the mean-field equation (1.3) in the large particle limit $J\to\infty$ was first proved rigorously in a recent paper [CBO-mfl-Huang2021] using a compactness argument, as we discuss further in Subsection 1.3. However, the method of proof employed in [CBO-mfl-Huang2021] does not provide an explicit rate of convergence with respect to the number of particles. Obtaining such an explicit rate of convergence is highly desirable, as it would enable to transfer convergence results concerning the mean-field equation (1.3) to the finite ensemble size setting (1.1) by means of a triangle inequality argument, modulo appropriate interfacing of the metrics. This is the approach proposed in [fornasier2021consensusbased, Theorem 13], where the authors obtain a partial mean-field result in a set of high probability, enabling them to prove convergence with high probability to the global minimizer in the finite ensemble size setting. Here, we provide a stronger version of the mean-field limit result given in [fornasier2021consensusbased, Proposition 15], extending to more general diffusion terms and providing scope for global convergence results in a stronger metric than convergence in probability.
@@
A number of extensions of CBO have been proposed in the literature. As mentioned previously, using an anisotropic diffusion term results in a scheme which is suitable for high-dimensional state spaces, and has been applied successfully in the training of artificial neural networks [CBO-high-dim]. Using componentwise geometric Brownian motion in this way allows to prove consensus formation directly at the particle level, which was achieved in [Ha2019, Ha2020ConvergenceAE, HA_Discrete_CBO_hetero_noise]. (On this subject, we point out that the results in [Ha2020ConvergenceAE] on the proximity of the consensus point to the global minimizer concern only the best case scenario; a more general convergence result in probability or a different metric is still lacking.) We also mention adaptations with mini-batching [jin2018random], memory effects [MR4160256], and Lévy noise [MR4553241], as well as extensions to multi-objective optimization [klamroth2022consensusbased, borghi2022adaptive] and global optimization with constraints [CBO-hypersurface-Fornasier_2020, MR4329816, CTV21, MR4538901, MR4404221, MR4337731]. Moreover, a variant tailored for objective functions with several minima was proposed in [bungert], and results establishing a link between CBO and PSO in the so-called small inertia limit were obtained in [CBO-PSO-zero-inertia-1, CBO-PSO-zero-inertia-2]. Many of the recent advances concerning the mathematical analysis of CBO and its extensions were recently overviewed in [CBO-survey-Totzeck2022]. Furthermore, a free software implementation of CBO and some of its extensions was recently published [Bailo2024].
@@ 1.2 Consensus-Based Sampling (CBS)
Consensus-Based Sampling is a sampling method inspired by the Ensemble Kalman Sampler (EKS) [EKS-1-Hofmann-Stuart-Li-GI-2019] and by consensus-based approaches such as CBO. It is based on an interacting particle system with a noise term preconditioned by the square root of the empirical covariance. The method in the finite ensemble size and continuous time setting reads
$\mathrm{d}X^{(j)}_{t}=-\Bigl{(}X^{(j)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\,\mathrm{d}W^{(j)}_{t},\qquad j\in\llbracket 1,J\rrbracket,$ (1.4)
@@
More precisely, the authors of [CBS-Carrillo2021] show that for uniformly convex $f$ and Gaussian initial data $\overline{\rho}_{0}$, the mean-field distribution $\overline{\rho}_{t}$ converges exponentially in Wasserstein distance as $t\to\infty$ to a Gaussian $\mathcal{N}(m_{\infty},C_{\infty})$ that can be made arbitrarily close to the Laplace approximation of $\rho_{f}$ by choosing a sufficiently large parameter $\beta$. In addition, it was noticed in [CBS-Carrillo2021] that, with $\lambda:=1$, the dynamics (1.4) behaves as a global optimization method for finding the global minimum of $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, just like the Consensus-Based Optimization method proposed in [CBO]. Since the convergence analysis in [CBS-Carrillo2021] is confined to the mean-field level, it is of great interest to prove rigorously that the particle dynamics (1.4) converges in an appropriate sense to the corresponding mean-field dynamics (1.5) as $J\to\infty$. This has so far been left as an open problem. Recently, consensus-based sampling was extended to enable sampling of multi-modal distributions in [bungert], and for rare event estimation in [2023arXiv230409077A].
@@ 1.3 Propagation of chaos
Although the consensus-based dynamics can be written in the form (1.6), it is not possible to do so in such a way that $\widetilde{b},\widetilde{\sigma},K_{b}$ and $K_{\sigma}$ are all globally Lipschitz continuous and at the same time $K_{b}$ and $K_{\sigma}$ are globally bounded. Therefore, the classical McKean’s theorem [ReviewChaintronII, Theorem 3.1] is not directly applicable, and a different approach is needed. We briefly discuss hereafter the existing mean-field results for CBO, noting that no results have been obtained so far for CBS.
  • In [CBO-mfl-Huang2021], propagation of chaos for CBO is proved through a compactness argument based on Prokhorov’s theorem. However, this method comes without convergence rates in terms of the number of particles $J$.
  • Later, propagation of chaos by means of synchronous coupling was shown with optimal rates in [CBO-hypersurface-Fornasier_2020] for a variant of CBO constrained to hypersurfaces.
  • Recent results in [fornasier2021consensusbased] also show a probabilistic mean-field approximation of the form
@@
    It is additionally proved in [fornasier2021consensusbased], by application of Markov’s inequality, that the set of realizations $\Omega_{M}$ has probability close to 0 for large $M$. Specifically, it holds that $\mathbf{P}(\Omega_{M})=\mathcal{O}(M^{-1})$.
  • Shortly before completing this work, we found out that propagation of chaos for CBO had been proved using a coupling approach very recently in [MR4553241]. In that work, the lack of global Lipschitz continuity of the weighted mean is handled by using stopping times, similarly to the approach we present in Subsection 5.1. There are, however, several differences between our results and those presented in [MR4553241]. Apart from the assumptions, which are slightly less restrictive in our work, the main technical differences are that (i) we improve existing Wasserstein stability estimates on the weighted mean (Corollary 3.3), (ii) we rely on a powerful result from the statistics literature [MR3696003, MR2597592] to bound one of the terms in Sznitman’s argument (Lemma 3.7) and (iii) we use a non-asymptotic central limit result to bound the probability that empirical measures exit a compact set (Subsection 2.3). Combined together, these ingredients enable to obtain mean-field result with the usual algebraic rate of convergence with respect to $J$, whereas the error scales as $\ln(\ln(J))^{-1}$ in the mean-field result obtained in [MR4553241].
@@ 1.4 Our contributions
  • We present two approaches for dealing with the weighted mean (in the case of CBO) and the weighted covariance (in the case of CBS). One of these approaches is based on the careful use of indicator sets, whereas the other is based on stopping times. The latter approach is inspired by a method initially proposed in [MR1949404] in order to handle local Lipschitz continuity issues in the context of numerical analysis for the Euler–Maruyama method.
@@ 2.1 Main assumptions
    for some constants $L_{f}>0$ and $s\geqslant 0$.
@@
    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
@@
  • In the particular case where $s=1$ and either $u=\ell=0$ or $u=\ell=2$, the assumptions in Assumption H coincide with those in [carrillo2018analytical]. The well-posedness and mean-field results proved in this paper generalize beyond this setting, as explained below.